OpenStudy (anonymous):

Combine the like terms to simplify the expression: 5 + 7 + 15xy2 + 9y3 + 10x –13xy2 – 3x – 8y3

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

Do you know what like terms are?

OpenStudy (anonymous):

Yes :P

OpenStudy (igreen):

Then simplify them..all you do is add or subtract them.

OpenStudy (igreen):

First list all of the like terms.

OpenStudy (anonymous):

There is exponents....

OpenStudy (igreen):

\(5 + 7 + 15xy^2 + 9y^3 + 10x –13xy^2 – 3x – 8y^3\) Like term pairs: 5 + 7 15xy^2 - 13xy^2 9y^3 - 8y^3 10x - 3x

OpenStudy (igreen):

Simplify all 4 of those.

OpenStudy (igreen):

Can you do that? @Raxxer

OpenStudy (igreen):

5 + 7 = 12 15xy^2 - 13xy^2 = 2xy^2 9y^3 - 8y^3 = y^3 10x - 3x = 7x

OpenStudy (igreen):

So we have: 12 + 2xy^2 + y^3 + 7x. Do you know how to put these in the correct order?

OpenStudy (igreen):

When you put it in order, you put in order of the greatest exponent. 12 = No exponent 2xy^2 = Exponent of 2 y^3 = Exponent of 3 7x = Exponent of 1 So the correct order is: y^3 + 2xy^2 + 7x + 12